The equation of motion of a particle is s =((t^3)-3t) where s is in metres and t is in
seconds. Find
i) the velocity and acceleration as functions of t,
ii) the acceleration after 2 seconds,
iii) the acceleration, when the velocity is 0.
1
Expert's answer
2020-09-30T19:33:31-0400

Consider the equation of motion of a particle "s=t^3-3t"


i)


Differentiate "s" with respect to "t" to obtain velocity function as,


"v=\frac{ds}{dt}"


="\frac{d}{dt}(t^3-3t)"


"=\frac{d}{dt}(t^3)-3\frac{d}{dt}(t)"


"=3t^2-3(1)"


"=3t^2-3"


Therefore, the velocity function is "v=3t^2-3"


Now, differentiate velocity function with respect to "t" to obtain acceleration function as,


"a=\frac{dv}{dt}"


"=\frac{d}{dt}(3t^2-3)"


"=3\frac{d}{dt}(t^2)-\frac{d}{dt}(3)"


"=3(2t)-0"


"=6t"


Therefore, the acceleration function is "a=6t"


ii)


Plug "t=2" into acceleration function to obtain acceleration after "2" seconds as,


"a=6(2)"


"=12"


Therefore, the acceleration after "2" seconds is "a=12" "m\/s^2"


iii)


When velocity is , then find the time as,


"3t^2-3=0"


"3(t^2-1)=0"


"t^2-1=0"


"t=1"


Here, negative value of "t" is neglected.


Now, plug "t=1" into acceleration function to obtain acceleration after "1" seconds as,


"a=6(1)"


"=6"


Therefore, the acceleration after "1" seconds is "a=6" "m\/s^2"